OpenStudy (anonymous):

Verify that the functions f and g are inverses of each other by showing that f(g(x))=x and g(f(x))=x. Please show all work. f(x)=ln(x-1) g(x)=1+e^x

OpenStudy (anonymous):

Simply: \[f(g(x))=\ln((1+e^x)-1)=x ???; g(f(x))=1+e^{\ln(x-1)}=x??\]

OpenStudy (anonymous):

You tell me :P

OpenStudy (anonymous):

So whats the answer?

OpenStudy (anonymous):

You tell me if f(g(x)) and g(f(x)) both equal x. I gave you the equations to solve.

OpenStudy (anonymous):

Idk this stuff thats why I need an answer .